Trace the curve y^2=(x+1)(x-1)^2 by showing all the properties you use to trace it?
1
Expert's answer
2019-09-11T09:28:13-0400
"f(x)=(x+1)(x-1)^2""f(x) \geqslant 0, x \geqslant -1; \,\, f(x) \geqslant 0, x \geqslant -1""f(x)=0 \Rightarrow x=-1, \,\, x=1""f'(x)=3x^2-2x-1 \Rightarrow x=-1/3, \,\, x=1""f \uparrow, x<-1/3, x>1; \,\, f \downarrow -1/3<x<1""y^2=f(x) \Rightarrow y=\pm \sqrt{f(x)}"

"\sqrt{f(x)}" have the same zeros as "f(x)" and it determining for "x \geqslant -1". Moreover, "y'=-\frac{f'(x)}{2\sqrt{f(x)}}" so we have "y \uparrow, -1<x<-1/3, x>1; \,\, y \downarrow -1/3<x<1". And we see that at "x=-1: \,\, y' \to \infty", so we have vertical tangent. And at "x \to 1 \pm: \,\,y'= \pm \sqrt{2}", i.e. parts of function have "90^{\circ}" angle between each other and "45^{\circ}" and "135^{\circ}" between "x"-axis (from right to left).

"y=-\sqrt{f(x)}" is given by symmetric mapping of "y=\sqrt{f(x)}" with respect to "x"-axis.